Mr. Rivera invest a part of P85,000.00 in a mortgage which pays 6% interest; and the rest in bonds paying 5% interest. The annual income from the bonds is twice the income from the mortgage. How much is invested in each of the 2 kinds of investment.
Answer #1

Mr. Rivera invest = 85,000

Lets say X from 85,000 is invested into mortgage which pays interest of 6% while (85,000 - X) is invested in bonds paying 5% interest.

Future value of mortgage after one years = X * 1.06 = 1.06X. Income from mortgage is 1.06X - X = 0.06X

Future value of bonds after one years = (85,000 - X) * 1.05 = 89,250 - 1.05X. Income from bonds is 89,250 - 1.05X - 85,000 + X = 4,250 - 0.05X

Annual income from bonds is double from the mortgage, thus 4,250 - 0.05X = 2 * 0.06X

0.17X = 4,250

X = 25,000

Thus, money invested in mortgage is 1,500 while invested in bonds is 3,000